The tollowing direct materials and direct labor data pertain to the operations of Laurel Company tor the month of August Costs Actual labor rate Actual materials price Standard labor rate Standard materials price $16 per hour $260 per ton $15.50 per hour $263 per ton uantities Actual hours incurred and used Actual quantity of materials purchased and used Standard hours used Standard quantity of materials used 3,700 hours 1,700 tons 3,780 hours 1,690 tons Compute the total, price, and quantity variances for materials and labor Total materials variance Materials price variance Materials quantity variance Total labor variance Labor price variance Labor quantity variance
Answer #1
Direct labor price variance = (SR – AR) x AH (15.5-16)*3700 = -1850
Direct labor quantity variance = SR x (SH – AH) 15.5*(3780-3700) = 1240
Total direct labor variance = (SR x SH) – (AR x AH) (3780*15.5)-(3700*16) = -610
=
Direct materials price variance = (SP – AP) x AQ (263-260)*1700 = 5100
Direct materials quantity variance = SP x (SQ – AQ) 263*(1690-1700) = -2630
Total direct materials variance = (SP x SQ) – (AP x AQ) (1690*263)-(260*1700) = 2470
where AR is Actual labour Rate
SR is Standard Labour Rate
AH is actual hours
SH is Standard hours
SP is Standard Price
AP is Actual Price
SQ is Standard Quantity
AQ is Actual Quantity